OpenStudy (anonymous):

Keisha is solving the equation 7^x = 9. Which shows the first step she should take? log 7^x = 9 7^x = log 9 log 7^x = ln 9 log 7^x = log 9

OpenStudy (anonymous):

First one.

OpenStudy (anonymous):

How do you know? Show me how you would work it out

OpenStudy (anonymous):

@hartnn @whpalmer4 Can you guys help me?

OpenStudy (anonymous):

Okay, hold on.

OpenStudy (campbell_st):

take the log of both sides of the equation

OpenStudy (campbell_st):

and make sure its the same base log.... so D

OpenStudy (anonymous):

Log should be placed first because it's conversion . 7^x = 9 This cannot be simplified without log log 7^x = 9....

OpenStudy (campbell_st):

the 1st line is \[\log(7^x) = \log(9)\] next apply the log law for powers \[x \times \log(7) = \log(9)\] now you can solve for x

OpenStudy (anonymous):

log 7^9 = x log 7/ log 9 = 1.9459101490553132/2.1972245773362196 = 0.8856218745807111 = 0.89

OpenStudy (anonymous):

@campbell_st The bases are NOT the same so it cannot be solved that way. 7 = 9 is not true so DO NOT put log on both sides.

OpenStudy (campbell_st):

well here is a simple idea... looking at the solution @YISODUMB15 has posted... test his solution given his answer of 0.89 does \[7^{0.89} = 9\] use a calculator to check

OpenStudy (whpalmer4):

No need for a calculator — raising 7 to a power < 1 isn't going to give you a number > 7!

OpenStudy (whpalmer4):

However, solving as @campbell_st suggests does give you the right answer...

OpenStudy (anonymous):

Oops

OpenStudy (anonymous):

Posted it again lol

OpenStudy (anonymous):

@whpalmer4 Exactly, Someone with sense. I solved it incorrectly but it isn't asking for a solution.

OpenStudy (whpalmer4):

@campbell_st has plenty of sense, and is adding some value by showing how you would continue on to the solution.

OpenStudy (campbell_st):

its asking for the 1st line of the solution... so you can get the correct solution

OpenStudy (anonymous):

The way he or she said it would be correct if it was 7^x = 7. However the bases are 7 and 9 which are not the same number. Also, I know @campbell_st has sense, her method is very useful, it just shouldn't be implied in this question.

OpenStudy (campbell_st):

the question is asking.... 7 to what power is equal to 9... 7 is the base, x is the power or exponent and 9 is the basic numeral...